If you push with a horizontal 650-N force to a 90-kg crate initially at resting on the floor, what must acceleration be? The static and kinetic energy coefficient of friction are us- 0.4, uk= 0.25 respectively.
Answer #1

Solution: F = 650 N m=90 kg 650-(0.4x90x9.8) 90a 650-352.8 = 90a 297.2-90a 297.2 90 a= 3.30 m/s EE 3.30 a= 2